OpenStudy (anonymous):

can someone please help me solve the inequality of -5x+8 <= x-4

OpenStudy (anonymous):

add 5x to both sides add 4 to both sides divide both sides by 6

OpenStudy (anonymous):

you want me to write it out?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[-5x+8\leq x-4\] add \(5x\) and get \[8\leq 6x-4\] add \(4\) to get \[12\leq 6x\] divide by 6 and get \[2\leq x\]

OpenStudy (anonymous):

Thank you so much

OpenStudy (anonymous):

yw